# The Conjugate Function Method for Surfaces with Elaborate Topological Types

### v1

@@ Abstract
The conjugate function method is an algorithm for numerical computation of conformal mappings for simply and multiply connected domains on surfaces. In this paper the conjugate function method is generalized and refined to achieve the same level of accuracy on simply and multiply connected planar domains and Riemann surfaces. The main challenge addressed here is the accurate and efficient construction of the conjugate problem for multiply connected domains. The method relies on high-order finite element methods which allow for highly accurate computations of mappings on surfaces, including domains of complex boundary geometry containing strong singularities and cusps. The efficacy of the proposed method is illustrated via an extensive set of numerical experiments with error estimates.
@@ 1 Introduction
Conformal geometry has many applications such as engineering (e.g. [5]) and mathematical geodesy (see [27]). Traditionally, these applications have been mostly restricted on conformal mappings between planar domains. Even in this setting finding mappings in the case of multiply connected domains has remained challenging. M.M.S. Nasser has used the boundary integral equation (BIE) with impressive results [19, 21]. Moreover, in a very recent work, Nasser discusses how to implement the generalized Koebe’s iterative method efficiently [15].
Koebe’s iterative method can also be implemented for surfaces. Gu, Luo, and Yau give an excellent overview of the state of the art in [6]. Due to both theoretical and implementation complexity, there are relatively few publications addressing specifics. See, for instance, [16, 17, 18, 25, 28]. Even though a properly implemented Koebe’s iteration converges reasonably fast (at least in the planar case), it is very difficult to provide any reliable error estimates or convergence rates.
Our method of choice for conformal mappings, the conjugate function method, was first introduced in [8] for simply and doubly connected planar domains. We use the $hp$-adaptive finite element method ($hp$-FEM) introduced in [11] for computing potential functions and conformal moduli, but other methods can be used as well, e.g., [13] the stochastic version of the algorithm was given for circular arc domains in the plane. One advantage of $hp$-FEM is that in error analysis than not only follows from the adoption of the finite element method for implementation but is also inherent in the form of the reciprocal identity relating the conformal capacities of the primary and conjugate problems. In this work this concept is made precise for surfaces in the specific setting of Q-type canonical domains.
On method development the main contribution of this work is the new, direct algorithm for the construction of the conjugate problem for multiply connected problems. Given the finite element discretization of the primary problem, the problem of finding the right boundary conditions for the conjugate problem, that is, the necessary Dirichlet-Neumann swap of the boundary conditions, can be reduced to a quadratic minimization problem that can be solved with a linear system of equations whose dimension is the number of the holes. Not only is the approach much faster than the previous ones, but the accuracy of the solution of the conjugate problem is brought to the same level as that of the primary problem. The same algorithm applies to problems on surfaces as well without any modifications. The conjugate function method is now comparable to the state-of-the-art methods with the additional benefits of the high-order finite element frameworks such as advanced error estimation and adaptivity.
@@ 1.1 Illustrative Example: 3D Face Recognition
Face recognition has many applications, especially in biometrics. Since a standard 2D image recognition is fragile due to many factors, 3D methods have been proposed as a more reliable alternative. Each 3D surface with suitable topology can be mapped conformally to a 2D domain and thus, the image recognition becomes a conformal map comparison problem. See [26].
In Figure 1 one example of how such a process might work is illustrated. First, four points on the outer boundary are fixed. In our formulation the canonical domain is a quadrilateral with horizontal slits if holes are present. Second, the Laplace-Beltrami equation is solved twice with different boundary conditions. As mentioned above, this process is only marginally slower if multiple holes are present. In the example both eyes and the mouth are modelled as holes.
@@ 1.2 Organization
The rest of the paper is organized as follows: After the introduction, preliminaries including the finite element method used and the planar conjugate function method are covered.
@@ 2 Preliminaries
In section some basic concepts and definitions used throughout this paper are introduced.
@@ 2.1 Riemann surfaces
Recall that a Hausdorff space $X$ is locally Euclidean of dimension $n\geq 1$ if every point $p$ in $X$ has a neighbourhood $U$ such that there is a homeomorphism $\varphi$ from $U$ onto an open subset of $\mathbb{R}^{n}$. We call the pair $(U,\ \varphi:U\to\mathbb{R}^{n})$ a chart, $U$ a coordinate neighbourhood or a coordinate open set, and $\phi$ a coordinate mapping or a coordinate system on $U$. We say that a chart $(U,\varphi)$ is a centered at $p\in U$ if $\varphi(p)=0$. Let $X$ be a two-dimensional manifold, which we also consider as a one-dimensional complex manifold. A complex chart on $X$ is a homeomorphisms $\varphi:U\to V$ from an open subset $U\subset X$ to an open subset $V\subset\mathbb{C}$. Two complex charts $\varphi_{j}:U_{j}\to V_{j}$ (for $j=1,2$) are said to be holomorphically compatible if the mapping
@@
The genus of a surface is defined as the largest number of non-intersecting simple closed curves that can be drawn on the surfaces without separating it. Intuitively, this corresponds to the number of holes on the surface. Note that in this paper we consider both closed surfaces and surfaces with multiple boundary components.
For more information of Riemann surfaces, we refer to [4].
@@ 2.2 High-Order Finite Element Method
High-order finite element methods have the capability for exponential convergence provided the discretization is constructed properly in both domain and in local polynomial order ($hp$-version). In many cases one has to work with cases where exact parameterization of the boundaries on the parameter space is not known. Then only algebraic convergence rate is to be expected.
@@ 2.2.1 Error Estimation
The solution method specific a posteriori error estimation method applied here is the so-called auxiliary subspace error estimation. The estimate measures the projection of the residual to the auxiliary space, that is, to a set of finite element degrees of freedom that extend the approximation space in a natural way. One example of such extension is to consider all edge modes of degree $p+1$ and bubble modes of $p+1$ and $p+2$, where $p$ is the (constant) polynomial degree used in the discretization. For details, see [7].
@@ 2.2.2 Special Types of Boundaries
One non-standard boundary type in the context of finite element method considered here is the slit or cut. We define a cut to be a set of element nodes and edges that form a loop with zero area. That is, the data structures support having two edges on top of each other, both with their own kinematic constraints.
@@ 3 The Conjugate Function Method On Planar Domains and Riemann Surfaces
Let $\Omega\subset\mathcal{S}$ is a simply connected domain on a Riemann surface $\mathcal{S}$ so that its boundary of $\Omega$ is a Jordan curve. We call $\Omega$ together with four positively oriented points $z_{1},z_{2},z_{3},z_{4}\in\partial\Omega$ a (generalized) quadrilateral and denote it by $Q=(\Omega;z_{1},z_{2},z_{3},z_{4})$. The boundary segments connecting the pairs of points $(z_{j},z_{j+1})$ for $j=1,2,3$, and $(z_{4},z_{1})$ for $j=4$, respectively, are denoted by $\gamma_{j}$.
@@
where $n$ is the unit exterior boundary normal and $\Delta_{S}$ is the Laplace-Beltrami operator. Then the conformal modulus is connected to the above boundary value problem by the identity (see e.g. Ahlfors [1, Theorem 4.5] and Papamichael and Stylianopoulos [22, Theorem 2.3.3]):
@@
For a quadrilateral $Q=(\Omega;z_{1},z_{2},z_{3},z_{4})$ we call $\widetilde{Q}=(\Omega;z_{2},z_{3},z_{4},z_{1})$ its conjugate quadrilateral and the corresponding problem (1) for quadrilateral $\widetilde{Q}$ the conjugate Dirichlet–Neumann problem. It is well-known that if $M(Q)=h>0$, then $M(\widetilde{Q})=1/h$, which leads to the following very useful reciprocal identity: For all quadrilaterals $Q$, we have
(3) $M(Q)M(\widetilde{Q})=1.$
Furthermore, we may observe that canonical conformal mapping of a quadrilateral $Q=(\Omega;z_{1},z_{2},z_{3},z_{4})$ onto the rectangle $R_{h}$ with vertices at $1+ih$, ihih, 0, and $1$, can be obtained by solving the corresponding Dirichlet–Neumann problem $u$ and its conjugate problem.
@@
Let $Q$ be a quadrilateral with modulus $h$, and suppose $u$ solves the Dirichlet–Neumann problem (1). If $v$ is a harmonic function conjugate to $u$, satisfying $v(\operatorname{Re}z_{3},\operatorname{Im}z_{3})=0$, and $\tilde{u}$ represents the harmonic solution for the conjugate quadrilateral $\widetilde{Q}$, then $v=h\tilde{u}$.
@@ 3.1 Q-type canonical domains
Let $h>0$, and let $\big{(}(\zeta_{1},d_{1}),\ldots,(\eta_{N},d_{N})\big{)}$ be complex numbers such that $\operatorname{Re}\zeta_{i}\in(0,1)$ and $\operatorname{Im}\zeta_{i}\in(0,h)$ for all $i=1,\ldots,N$, where $N\geq 1$. For a multiply connected plane case, we will consider canonical domains of the type
(4) $\big{(}(0,1)\times(0,h)\big{)}\setminus\bigcup_{j=1}^{N}[\zeta_{j},\zeta_{j}+d_{j}],$
@@ 4.1 Reciprocal Identity for Multiply Connected Domains
Let $\Omega$ be a domain on an oriented surface $S$ that can mapped conformally onto a Q-type canonical domain $Q_{*}=(\Omega;z_{1},z_{2},z_{3},z_{4})$ defined by (4), where the points $z_{1},z_{2},z_{3},z_{4}$ are chosen from the same boundary component of $\Omega$. Then the generalized conformal moduli of $\Omega$ satisfy the reciprocal identity:
@@ 4.2 Linear Algebra Based Construction of the Conjugate Problem
Let us divide the degrees of freedom, internal and boundary, into sets. The sets of a discretized quadrilateral are denoted $B,D_{0},D_{1},N^{0}$, and $N^{1}$, for internal, Dirichlet $u=0$, Dirichlet $u=1$, Neumann with Dirichlet $u=0$ in the primary problem, and Neumann with Dirichlet $u=1$ in the primary problem, degrees of freedom, respectively. The degrees of freedom on the boundaries of the holes $E_{i}$ are all Dirichlet, constant over the boundary and initially unknown $\delta_{E_{i}}$.
The discretized system is
@@
In order to derive the quadratic form we first divide the degrees of freedom into free, that is, zero Neumann, and fixed, where the internal and Neumann degrees of freedom are free, and all Dirichlet ones are fixed. These sets are denoted by $I$ and $D$, respectively. The matrix $A$ has a simple $2\times 2$ form $A=\left(\begin{smallmatrix}A_{II}&A_{ID}\\ A_{DI}&A_{DD}\\ \end{smallmatrix}\right)$, and $x=\left(\begin{smallmatrix}x_{I}&x_{D}\end{smallmatrix}\right)^{T}$. Using the block form one can solve (formally) the free degrees of freedom
@@
Using this, one can then evaluate the quadratic form
@@
Let us next denote the vector of unknown parameter values $x^{\prime}_{D}=\left(\begin{smallmatrix}\delta_{E_{1}}&\dots&\delta_{E_{n}}\end{smallmatrix}\right)^{T}.$ Then we arrive at the quadratic minimization problem
@@
where $K$ and $b$ are the coefficient matrix (symmetric) and vector, respectively, extracted from (12). We conclude that the problem of finding the Dirichlet boundary values of the holes reduces to finding a solution to a $n\times n$ linear system of equations. Of course, this construction requires that the necessary bookkeeping of degrees of freedom is available.
@@ 4.3 Details of the Reduction Step
The idea is to construct a reduction matrix $R$ which maps, in other words, sums, all coefficients of the dofs $x_{E_{j}}$ to those of $\delta_{E_{j}}$. The matrix $R\in\mathbb{R}^{m\times(n+1)}$ has the form
@@ 4.4 Effect of the Improved Construction: Two Holes on a Sphere Revisited
The precise description of the problem on the parameter plane is as follows: The computational domain is a unit circle centred at $(1/2,1/2)$, with two circular holes with radius =1/4=1/4 and centres at $(1/4,1/4)$ and $(3/4,3/4)$ denoted as $B_{1}$ and $B_{2}$, respectively. The four corners are selected symmetrically at $z_{i}=(1/2+\cos(\pi(k-1)/2),1/2+\sin(\pi(k-1)/2))$, $k=1,\dots,4$. The optimization step gives us potentials $v_{1}=0.5344459257688663$ on $\partial B_{1}$ and $v_{2}=0.4883797948571419$ on $\partial B_{2}$. The moduli obtained with the new construction of the conjugate problem are
(14) $M(Q)=0.7901907691109106,\;M(\widetilde{Q})=1.2655172762842442,$
with $\operatorname{reci}(Q)=6.987019163240404\times 10^{-8}$.
Figure 2: Two Holes Revisited. Convergence of the reciprocal error as a function of degrees of freedom using optimization and the new construction. The current version is eventually more accurate on exactly the same mesh, and dips below the threshold limit of $10^{-7}$ that indicates the best possible accuracy in the previous version (loglog-plot).

### v2

@@ Abstract
The conjugate function method is an algorithm for numerical computation of conformal mappings for simply and multiply connected domains on surfaces. In this paper the conjugate function method, earlier used for simply connected domains, is generalized and refined to achieve the same level of accuracy on multiply connected planar domains and Riemann surfaces. The main challenge is the accurate and efficient construction of boundary values for the conjugate problem on multiply connected domains. The method relies on high-order finite element methods which allow for highly accurate computations of mappings on surfaces, including domains of complex boundary geometry containing strong singularities and cusps. We also derive the reciprocal error estimate for the multiply connected case. The efficacy of the proposed method is illustrated via an extensive set of numerical experiments with error estimates.
@@ 1 Introduction
Conformal geometry has many applications such as engineering (e.g. [5]) and mathematical geodesy (see [27]). Traditionally, numerical methods of approximating conformal mappings have been mostly restricted to planar domains. Even in this setting finding mappings between multiply connected domains has remained challenging. M.M.S. Nasser has used the boundary integral equation (BIE) with impressive results [19, 21]. Furthermore, in a very recent work, Nasser discusses efficient implementation of the generalized Koebe’s iterative method [15].
Koebe’s iterative method can also be implemented for surfaces. Gu, Luo, and Yau give an excellent overview of the state of the art in [6]. Due to both theoretical and implementation complexity, there are relatively few publications addressing specifics. See, for example, [16, 17, 18, 25, 29]. Even though a properly implemented Koebe’s iteration converges reasonably fast (at least in the planar case), it is very difficult to provide any reliable error estimates or convergence rates.
Our method of choice for conformal mappings, the conjugate function method, was first introduced in [8] for simply and doubly connected planar domains. We use the $hp$-adaptive finite element method ($hp$-FEM) introduced in [11] to compute potential functions and conformal moduli, but other methods can also be used, e.g., in [13] the stochastic version of the conjugate function algorithm was given for circular arc domains in the plane. An advantage of $hp$-FEM is the mature error analysis. Interestingly, the reciprocal identity relating the conformal capacities of the primary and conjugate problems can also be viewed as an error indicator. In this work this concept is made precise for surfaces with multiple boundary components in the specific setting of so-called Q-type canonical domains (see [9] and below).
On method development the main contribution of this work is the new, direct algorithm for the construction of the conjugate problem for multiply connected problems. Given the finite element discretization of the primary problem, the problem of finding the right boundary conditions for the conjugate problem, that is, the necessary Dirichlet-Neumann swap of the boundary conditions, can be reduced to a quadratic minimization problem that can be solved with a linear system of equations whose dimension is the number of the holes. Not only is the approach much faster than the previous ones, but the accuracy of the solution of the conjugate problem is brought to the same level as that of the primary problem. The same algorithm applies to problems on multiply connected domains on surfaces as well without any modifications.
@@ 1.1 Illustrative Example: 3D Face Recognition
Face recognition has many applications, especially in biometrics. Since a standard 2D image recognition is fragile due to many factors, 3D methods have been proposed as a more reliable alternative. Each 3D surface with suitable topology can be mapped conformally to a 2D domain, and thus, the image recognition becomes a conformal map comparison problem. See [26].
In Figure 1 one example of how such a process might work is illustrated. First, four points on the outer boundary are fixed. In our formulation, the canonical domain is a quadrilateral with horizontal slits if holes are present. Second, the Laplace-Beltrami equation is solved twice with different boundary conditions. As mentioned above, this process is only marginally slower if multiple holes are present. In the example, both eyes and the mouth are modeled as holes.
@@ 1.2 Organization
The rest of the paper is organized as follows: After the introduction, preliminaries including the finite element method used are presented. In Section 3 the conjugate function method is introduced. Section 4 discusses both the proof of the reciprocal identity for oriented surfaces, and the efficient construction of the conjugate problem. Torii are given as a special class of closed surfaces in Section 5. Sections 6-8 cover series of numerical experiments and applications, before conclusions at the end.
@@ 2 Preliminaries
In this section, some basic concepts and definitions used throughout this paper are introduced.
@@ 2.1 Riemann surfaces
Let $X$ be a Hausdorff space that is locally homeomorphic to the Euclidean plane $\mathbb{R}^{2}$, which we also consider as a one-dimensional complex manifold. We assume that its boundary $\Gamma$ is either empty or consists of a finite number of rectifiable arcs that may intersect only at their endpoints, and its closure $X\cup\Gamma$ is compact. A complex chart on $X$ is a homeomorphism $\varphi:U\to V$ from an open subset $U\subset X$ onto an open subset $V\subset\mathbb{C}$. Two complex charts $\varphi_{j}:U_{j}\to V_{j}$ (for $j=1,2$) are said to be holomorphically compatible if $U_{1}\cap U_{2}=\emptyset$ or the mapping
@@
The genus of a surface is defined as the largest number of non-intersecting simple closed curves that can be drawn on the surfaces without separating it. Intuitively, this corresponds to the number of holes on the surface. Note that in this paper, we consider both closed surfaces and surfaces with multiple boundary components.
For more information about Riemann surfaces, we refer to [4, 14].
@@ 2.2 High-Order Finite Element Method
High-order finite element methods have the capability for exponential convergence in conformal capacity problems provided that for a given discretization the polynomial order (hp-version) is properly chosen and the discretization is correctly adjusted to reflect the intricacies of the domain boundary. In many cases one has to work with cases where the exact parameterization of the boundaries on the parameter space is not known. Then only algebraic convergence rate is to be expected.
@@ 2.2.1 Error Estimation
The a posteriori error estimation method specific to the solution method applied here is the so-called auxiliary subspace error estimation. The estimate measures the projection of the residual to the auxiliary space, that is, to a set of finite element degrees of freedom that extend the approximation space in a natural way. One example of such extension is to consider all edge modes of degree $p+1$ and bubble modes of $p+1$ and $p+2$, where $p$ is the (constant) polynomial degree used in the discretization. For details, see [7].
@@ 2.2.2 Special Types of Boundaries
One non-standard boundary type in the context of finite element method considered here is the slit. We define a slit to be a set of element nodes and edges that form a loop with zero area. That is, the data structures support having two edges on top of each other, both with their own kinematic constraints.
@@ 3 The Conjugate Function Method On Planar Domains and Riemann Surfaces
Let $\Omega\subset\mathcal{S}$ be a simply connected domain on a Riemann surface $\mathcal{S}$ so that the boundary of $\Omega$ is a Jordan curve. We call $\Omega$ together with four positively oriented points $z_{1},z_{2},z_{3},z_{4}\in\partial\Omega$ a (generalized) quadrilateral and denote it by $Q=(\Omega;z_{1},z_{2},z_{3},z_{4})$. The boundary segments connecting the pairs of points $(z_{j},z_{j+1})$ for $j=1,2,3$, and $(z_{4},z_{1})$ for $j=4$, respectively, are denoted by $\gamma_{j}$.
@@
where $n$ is the unit exterior boundary normal and $\Delta_{S}$ is the Laplace-Beltrami operator. For $\Omega\subset\mathbb{C}$ the conformal modulus is connected to the above boundary value problem by the identity (see e.g. Ahlfors [1, Theorem 4.5] and Papamichael and Stylianopoulos [22, Theorem 2.3.3]):
@@
For a quadrilateral $Q=(\Omega;z_{1},z_{2},z_{3},z_{4})$ we call $\widetilde{Q}=(\Omega;z_{2},z_{3},z_{4},z_{1})$ its conjugate quadrilateral and the corresponding problem (1) for the quadrilateral $\widetilde{Q}$ the conjugate Dirichlet–Neumann problem. It is well-known that if $M(Q)=h>0$, then $M(\widetilde{Q})=1/h$, which leads to
(3) $M(Q)M(\widetilde{Q})=1,$
for all quadrilaterals $Q$. This reciprocal identity is very useful since it can be interpreted as an error estimate [11].
Furthermore, we may observe that the canonical conformal mapping of a quadrilateral $Q=(\Omega;z_{1},z_{2},z_{3},z_{4})$ onto the rectangle $R_{h}$ with vertices at $1+ih$, ihih, 0, and $1$, can be obtained by solving the corresponding Dirichlet–Neumann problem and its conjugate problem.
@@
Let $Q$ be a quadrilateral with modulus $h$, and suppose $u$ solves the Dirichlet–Neumann problem (1). If $v$ is a harmonic function conjugate to $u$, satisfying $v(\operatorname{Re}z_{3},\operatorname{Im}z_{3})=0$, and $\tilde{u}$ represents the harmonic function solving the Dirichlet–Neumann problem for the conjugate quadrilateral $\widetilde{Q}$, then $v=h\tilde{u}$.
@@ 3.1 Q-type canonical domains
Let $h>0$, and let $\big((\zeta_{1},d_{1}),\ldots,(\zeta_{N},d_{N})\big)$ be complex numbers such that $\operatorname{Re}\zeta_{i}\in(0,1)$ and $\operatorname{Im}\zeta_{i}\in(0,h)$ for all $i=1,\ldots,N$, where $N\geq 1$. For the case of multiply connected planar domains, we will consider canonical domains of the type
(4) $\big((0,1)\times(0,h)\big)\setminus\bigcup_{j=1}^{N}[\zeta_{j},\zeta_{j}+d_{j}]\subset\mathbb{C},$
@@ 4.1 Reciprocal Identity for Multiply Connected Domains
Let $\Omega$ be a domain on an oriented surface $S$ that can be mapped conformally onto a Q-type canonical domain $Q_{*}=(\Omega;z_{1},z_{2},z_{3},z_{4})$ defined by (4), where the points $z_{1},z_{2},z_{3},z_{4}$ are chosen from the same boundary component of $\Omega$. Then the generalized conformal moduli of $\Omega$ satisfy the reciprocal identity:
@@
Remark 4.3.
Previously, the use of reciprocal identity as a method of error estimation has been verified only in the simply connected case [8].
@@ 4.2 Linear Algebra Based Construction of the Conjugate Problem
As indicated in (8), the discretized primary and conjugate problems share the same degrees of freedom before application of boundary conditions. The degrees of freedom admit a partition that can be used to define a block structure for the discretized matrix $A$. For the conjugate problem the degrees of freedom on the four boundary segments $\gamma_{i}$ defining the quadrilateral $Q$ are denoted by $D_{0},D_{1},N^{0}$, and $N^{1}$, where $D_{0}$ indicates $v=0$ and $N^{0}$ zero Neumann on the segment with $u=0$ on the primary problem. The internal degrees of freedom are denoted by $B$. Finally, the degrees of freedom on the boundaries of the holes are denoted by $E_{i}$ .
Remark 4.4.
Since the potential is constant over the boundary of every hole in the conjugate problem, all edge degrees of freedom are set to zero. This is a technical detail that does not affect the discussion below.
Using the notation given above the discretized system has the following block structure
@@
In order to derive the quadratic form we further divide the degrees of freedom into two classes: free ($B$, $N^{0}$, $N^{1}$) and fixed (all Dirichlet). These classes are denoted by $I$ and $D$, respectively. Then the matrix $A$ has a simple $2\times 2$ form $A=\left(\begin{smallmatrix}A_{II}&A_{ID}\\ A_{DI}&A_{DD}\\ \end{smallmatrix}\right)$, and $x=\left(\begin{smallmatrix}x_{I}&x_{D}\end{smallmatrix}\right)^{T}$. Using the block form one can solve (formally) the free degrees of freedom
@@
Therefore, one can simplify the quadratic form
@@
Let us next denote the vector of unknown parameter values $x^{\prime}_{D}=\left(\begin{smallmatrix}v_{1}&\dots&v_{n}\end{smallmatrix}\right)^{T}.$ Using the fact that every nonzero degree of freedom on $E_{i}$ is constant and equal to $v_{i}$, we arrive at the quadratic minimization problem
@@
where $K$ and $b$ are the coefficient matrix (symmetric) and vector, respectively, extracted from (12). We conclude that the problem of finding the Dirichlet boundary values of the holes reduces to finding a solution to an $n\times n$ linear system of equations. Of course, this construction requires that the necessary bookkeeping of degrees of freedom is available.
@@ 4.3 Details of the Reduction Step
The idea is to construct a reduction matrix $R$ which maps, in other words, sums, all coefficients of the dofs $x_{E_{j}}$ to those of $v_{j}$. The matrix $R\in\mathbb{R}^{m\times(n+1)}$ has the form
@@ 4.4 Effect of the Improved Construction: Two Holes on a Sphere Revisited
The precise description of the problem on the parameter plane is as indicated in Table 1. The moduli obtained with the new construction of the conjugate problem are
(14) $M(Q)=0.7901908,\;M(\widetilde{Q})=1.2655173,$
with $\operatorname{reci}(Q)=7\times 10^{-8}$.
Table 1: Two Holes: Unit circle centered at $(1/2,1/2)$. The locations, radii, and the obtained potential values in the construction of the conjugate problem are given for both holes.
Hole Center Radius $v_{k}$ on $\partial B_{k}$
$B_{1}$ $(1/4,1/4)$ $1/4$ $0.5344459257688663$
$B_{2}$ $(3/4,3/4)$ $1/4$ $0.4883797948571419$
Figure 2: Two Holes Revisited. Convergence of the reciprocal error as a function of degrees of freedom using optimization and the new construction. The current version is eventually more accurate on exactly the same mesh, and dips below the threshold limit of $10^{-7}$ that indicates the best possible accuracy when the black-box optimization routine was used (loglog-plot) [10].